# Confidence Intervals for Linear Models with Arbitrary Noise Contamination

### v1

@@ 1 Introduction
In this paper, we propose an algorithm that computes a confidence interval for a given coordinate of $b$. Our proposed algorithm does not depend on the knowledge of $\epsilon$. We show that the confidence interval not only has the coverage property, but its length is also bounded by $O_{\mathbb{P}}\left(\frac{1}{\sqrt{n(1-\epsilon)^{2}}}\right)$. According to the lower bound for estimation in d’Orsi et al., (2021), this is the optimal rate even if the value of $\epsilon$ is known. In other words, unlike the negative result of Luo and Gao, (2024), there is no adaptation cost in robust confidence interval construction in the regression setting. The length of the confidence interval automatically shrinks as the quality of the data increases (i.e. a smaller value of $\epsilon$). Moreover, the same theoretical guarantee holds if the Gaussian noise $N(0,1)$ in (2) is replaced by other non-degenerate distributions such as $t$ or Cauchy. Thus, our confidence interval is robust against both heavy tailed noise and arbitrary contamination.
@@
It is interesting to note that (7) can also be derived from the general strategy of constructing adaptive robust confidence interval proposed by Luo and Gao, (2024). Specifically, as long as one can construct a testing function that is able to distinguish between different values of the parameter together with different contamination proportions, the testing fuction can be inverted into a robust confidence interval that is adaptive to the unknown contamination level $\epsilon$. Although they only stated this framework for location families, it turns out that the same principle can be applied to the regression setting. In particular, one can construct an appropriate test whose inversion recovers the formula (7). We refer the reader to Section 3.5 for more details.
@@ 1.1 Related Work
Due to the lack of assumption on contamination, confidence interval construction under the noise (2) is required to be nonparametric, since there is no asymptotic distribution available. One class of nonparametric methods is based on the boot bootstrap. In particular, bootstrap in linear regression was developed in Freedman, (1981), and its variants and theoretical properties were further investigated by Bickel and Freedman, (1983); Mammen, (1989, 1993). However, bootstrap consistency is usually based on the existence of an asymptotic distribution, which is still violated by (2).
@@ 1.2 Paper Organization
We introduce the new algorithm and state its theoretical guarantee in Section 2. To understand why the proposed method works, we will first discuss the problem and illustrate the main idea for $p=1$ in Section 3. Then in Section 4, we will discuss how to reduce the original problem to the one-dimensional setting through decorrelation. Extensive numerical experiments will be reported in Section 5. Finally, we collect all technical proofs in Section 6.
@@ 2 Main Results
In words, for each direction $v\in S_{p-1}$, the projection $v^{T}X_{i}$ is a symmetric sub-Gaussian random variable whose distribution is not degenerate in a neighborhood of zero.
@@
Consider i.i.d. samples $\{(x_{i},w_{i},y_{i})\}_{i=1}^{n}$ generated from the linear model (8), with $\{(x_{i},w_{i})\}_{i=1}^{n}$ and $\{(z_{i})\}_{i=1}^{n}$ independent from each other and satisfying Assumption A and (2). For any $\alpha\in(0,1)$, there exist constants $c>0$ and $C>0$ depending on $\alpha$, such that the interval $\widehat{\beta}_{L},\widehat{\beta}_{R}]$ computed by Algorithm 1 satisfies
@@
The noise condition (2) assumes that $z_{i}$’s are generated from a standard $N(0,1)$ if they are not drawn from contamination. Our proofs would remain the same if $N(0,1)$ is replaced by any distribution satisfying $\mathbb{P}(|z_{i}|\leq 1)\gtrsim 1$, which includes most heavy tailed distributions such as $t$ and Cauchy.
@@
Assumption A on the covariate assumes symmetry and sub-Gaussianity. The symmetry condition can be avoided by incorporating a preprocessing step of computing the pairwise difference $y_{i}-y_{j}=(X_{i}-X_{j})^{T}\beta+(z_{i}-z_{j})$. The new covariate $X_{i}-X_{j}$ certainly satisfies symmetry regardless of the distribution of $X_{i}$. The sub-Gaussianity condition is imposed for the convenience of certain technical arguments in the proof, and we do not think it is necessary for the result to hold. It can possibly be relaxed with some lengthier and more tedious arguments.
We finally give a brief discussion on the condition $\frac{p^{2}}{n(1-\epsilon)^{4}}\leq c$ for some sufficiently small constant $c>0$. When $\epsilon$ is bounded away from $1$, the condition allows $p$ to be a growing function of $n$ as long as $n\gg p^{2}$. The requirement that the sample size needs to exceed squared degrees of freedom has been previously noted in the high-dimensional literature when signal is very sparse (Zhang and Zhang,, 2014; Javanmard and Montanari,, 2014; van de Geer et al.,, 2014), and can be shown to be necessary for interval estimation with parametric rate (Ren et al.,, 2015; Cai and Guo,, 2017; Verzelen and Gassiat,, 2018). When the signal is dense, however, $n\gg p^{2}$ is actually not needed for confidence interval construction when the data has no contamination (Wen et al.,, 2025). Whether or not $\frac{p^{2}}{n(1-\epsilon)^{4}}\leq c$ can be further relaxed under Huber’s contamination model (2) is an interesting theoretical question to be explored in the future.
@@ 3.1 A Symmetric Location Model
where $z_{i}\sim(1-\epsilon)N(0,1)+\epsilon Q_{0}$ for some distribution $Q_{0}$ that is symmetric around zero. In other words, $z_{i}\sim Q_{0}$ implies $z_{i}\sim Q_{0}$. We emphasize that such symmetry assumption is not assumed for the error distribution in (9).
@@ 3.3 Application in Robust Regression
Since the covariate $x_{i}$ follows $N(0,1)$ and the noise variable $z_{i}$ is independent from $x_{i}$, the random variable $\frac{z_{i}}{x_{i}}$ is symmetric around zero regardless of the noise distribution. Therefore, (15) is an instance of the symmetric location model (9). Even if the distribution of $x_{i}$ is not necessarily $N(0,1)$ or not even symmetric, one could still deal with a linear model with symmetric design by preprocessing the data with pairwise difference $y_{i}-y_{j}=\beta(x_{i}-x_{j})+z_{i}-z_{j}$.
@@ 4.1 Decorrelating Nuisance Parameters
Even though $\widetilde{x}_{i}$ and $(\gamma-\widehat{\gamma})^{T}w_{i}+z_{i}$ are still not independent in (27), the independence between $\widetilde{x}_{i}$ and $z_{i}$ and the uncorrelatedness between $\widetilde{x}_{i}$ and $w_{i}$ are sufficient for us to apply (21). We emphasize that the uncorrelatedness condition (28) is only in the sample level, which turns out to be particularly convenient when bounding the model missspecification error. Moreover, the $\widehat{\gamma}$ in (26) a robust estimator for $\gamma=\beta\widehat{\alpha}+\theta$, which will ensure that the scale of the noise variable $(\gamma-\widehat{\gamma})^{T}w_{i}+z_{i}$ is well controlled.
@@ 4.2 Coverage
To analyze the effect of model missspecification caused by $(\gamma-\widehat{\gamma})^{T}w_{i}$, we use the smoothness of $g(\cdot)$ and have
@@ 5.1 Numerical Setups
    Gaussian: $z_{i}\overset{iid}{\sim}(1-\epsilon)N(0,1)+\epsilon\left(\frac{1}{2}N(10^{2},1)+\frac{1}{2}N(10^{4},1)\right)$.
@@
    Cauchy: $z_{i}\overset{iid}{\sim}(1-\epsilon)C(0,1)+\epsilon\left(\frac{1}{2}N(10^{2},1)+\frac{1}{2}N(10^{4},1)\right)$.
@@
The threshold level $1.5\sqrt{\log(2/\alpha)}$ used in Algorithm 1 is to guarantee non-asymptotic coverage in Theorem 1. One may wonder whether replacing $1.5\sqrt{\log(2/\alpha)}$ with the more aggressive Gaussian quantile $\Phi^{-1}(1-\alpha/2)$ works better in practice. While the Gaussian approximation works in the asymptotics of $\frac{p^{2}}{n(1-\epsilon)^{4}}\rightarrow 0$, our numerical results indicates that it should be used with caution when the sample size $n$ is small or $\epsilon$ is large. In addition to this variation of Algorithm 1, we also implement several methods in the literature. This includes the residual bootstrap (RB) of Freedman, (1981), the restructured regression (RR) of Zhu and Bradic, (2018), the cyclic permutation test (CPT) of Lei and Bickel, (2021), the convex hull method (HulC) of Kuchibhotla et al., (2024) and the residual permutation test (RPT) of Wen et al., (2025). We also report the performance of the classical confidence interval computed from t-statistic centering at the ordinary least-squares estimator (OLS). Among the above methods, RR (Zhu and Bradic,, 2018), CPT (Lei and Bickel,, 2021) and RPT (Wen et al.,, 2025) are originally proposed to test the null hypothesis that a coefficient is zero, which also lead to confidence intervals by inverting the tests.
@@ 5.2 Implementation Details
where $\rho(\cdot)$ is the same Huber loss used in Algorithm 1. The boostrap sample size is taken as $200$. The restructured regression method has two versions in Zhu and Bradic, (2018), and we implement the one with the knowledge of the design covariance. The cyclic permutation test (Lei and Bickel,, 2021) is implemented with the recommended preordering computed using a genetic algorithm. The genetic algorithm is computed with $1000$ random samples. The HulC method (Kuchibhotla et al.,, 2024) is also implemented with the Huber regression (45), together with a recommended randomized choice of sample splitting to prevent the interval from being overly conservative. The residual permutation test (Wen et al.,, 2025) is implemented with the more powerful version using the empirical p-values, and the tuning parameters are set as recommended in the original work.

### v2

@@ 1 Introduction
A natural approach for handling unknown contamination levels is Lepski’s method (Lepski et al.,, 1999; Jain et al.,, 2022), which has been successfully used for adaptive estimation. However, while such methods yield estimators that achieve optimal rates without knowing $\epsilon$, they do not directly provide confidence intervals with guaranteed coverage, as uncertainty quantification requires estimating the statistical error itself rather than merely selecting an estimator with optimal rate.
In this paper, we propose an algorithm that computes a confidence interval for a given coordinate of $b$. The procedure does not require knowledge of the contamination level $\epsilon$. We show that the resulting confidence interval achieves the desired coverage uniformly over all contamination distributions $Q$, and its length is bounded by $O_{\mathbb{P}}\left(\frac{1}{\sqrt{n(1-\epsilon)^{2}}}\right)$. In particular, this rate is minimax-optimal up to constants even when $\epsilon$ is known (d’Orsi et al.,, 2021). Taken together, these results establish that one can simultaneously attain adaptivity to unknown contamination, minimax-optimal length, and robustness over arbitrary contamination distributions. In contrast to the negative result of Luo and Gao, (2024), this shows that there is no adaptation cost for robust confidence interval construction in the regression setting. The length of the confidence interval automatically shrinks as the data quality improves (i.e., as $\epsilon$ decreases). Moreover, the same theoretical guarantee continues to hold when the Gaussian noise $N(0,1)$ in (2) is replaced by other non-degenerate distributions such as $t$-distribution or the Cauchy distribution, demonstrating robustness to both heavy-tailed noise and arbitrary contamination.
@@
It is interesting to note that (7) can also be derived from the general strategy of constructing adaptive robust confidence interval proposed by Luo and Gao, (2024). Specifically, as long as one can construct a testing function that is able to distinguish between different values of the parameter together with different contamination proportions, the testing function can be inverted into a robust confidence interval that is adaptive to the unknown contamination level $\epsilon$. Although they only stated this framework for location families, it turns out that the same principle can be applied to the regression setting. In particular, one can construct an appropriate test whose inversion recovers the formula (7). We refer the reader to Section 3.5 for more details.
@@ 1.1 Related Work
Due to the lack of assumptions on contamination, confidence interval construction under the noise (2) is required to be nonparametric, since there is no asymptotic distribution available. One class of nonparametric methods is based on the bootstrap. In particular, bootstrap in linear regression was developed in Freedman, (1981), and its variants and theoretical properties were further investigated by Bickel and Freedman, (1983); Mammen, (1989, 1993). However, bootstrap consistency is usually based on the existence of an asymptotic distribution, which is still violated by (2).
@@ 1.2 Paper Organization
We introduce the new algorithm and state its theoretical guarantee in Section 2. To understand why the proposed method works, we will first discuss the problem and illustrate the main idea for $p=1$ in Section 3. Then in Section 4, we will discuss how to reduce the original problem to the one-dimensional setting through decorrelation. Extensive numerical experiments will be reported in Section 5. We discuss two robust regression settings beyond the oblivious contamination model in Section 6 and collect all technical proofs in Section 7.
@@ 2 Main Results
In words, for each direction $v\in S_{p-1}$, the projection $v^{T}X_{i}$ is a symmetric sub-Gaussian random variable whose distribution is not degenerate in a neighborhood of zero. Note that the gradient of the Huber loss $\rho(t)$ is non-constant only when $|t|$ is small, which is a crucial property for robustness against large outliers. The non-degeneracy in a neighborhood of zero is thus necessary for regression with Huber loss.
@@
Consider i.i.d. samples $\{(x_{i},w_{i},y_{i})\}_{i=1}^{n}$ generated from the linear model (8), with $\{(x_{i},w_{i})\}_{i=1}^{n}$ and $\{z_{i}\}_{i=1}^{n}$ independent from each other and satisfying Assumption A and (2). For any $\alpha\in(0,1)$, there exist constants $c>0$ and $C>0$ depending on $\alpha$, such that the interval $\widehat{\beta}_{L},\widehat{\beta}_{R}]$ computed by Algorithm 1 satisfies
@@
The noise condition (2) assumes that $z_{i}$’s are generated from a standard $N(0,1)$ if they are not drawn from contamination. Our proofs would remain the same if $N(0,1)$ is replaced by any distribution satisfying $\mathbb{P}(|z_{i}|\leq 1)\gtrsim 1$, which includes most heavy tailed distributions such as $t$ and Cauchy. More generally, if the noise is scaled by $\sigma$ (e.g. $N(0,\sigma^{2})$), we can scale Algorithm 1 accordingly by $g(\cdot)\rightarrow g(\cdot/\tau)$ and $\rho(\cdot)\rightarrow\rho(\cdot/\tau)$ for some $\tau>0$. Then, Theorem 1 continues to hold as long as $\tau\lesssim\sigma$ and $\frac{p^{2}}{n(1-\epsilon)^{4}(\tau/\sigma)^{4}}$ is sufficiently small.
Remark 2.
The sub-Gaussianity of the design can be relaxed to weaker conditions, and the result would still hold under stronger conditions on dimension. For example, under the sub-Weibull tail ${\rm{exp}}\left(-(t/K_{1})^{\kappa}\right)$ for some $\kappa\in(0,2]$, the condition $\frac{p^{2}}{n(1-\epsilon)^{4}}\leq c$ would be generalized to $\frac{p^{1+\frac{2}{\kappa}}}{n(1-\epsilon)^{4}}\leq c$. The same proof technique still works after incorporating concentration inequalities in Kuchibhotla and Chakrabortty, (2022).
Remark 3.
It is crucial that the interval length guarantee holds with high probability instead of in expectation. It can be shown that $\mathbb{E}|\widehat{\beta}_{R}-\widehat{\beta}_{L}|=\infty$ (see Proposition 5) in the worst-case. This is because with probability $\epsilon^{n}$, all $z_{i}$’s are drawn from $Q$ under (2), and thus the expectation is at least $\epsilon^{n}$ times the expectation under $Q$, which can be arbitrarily large.
@@
Assumption A on the covariate assumes symmetry. The symmetry condition can be avoided by incorporating a preprocessing step of computing the pairwise difference $y_{i}-y_{j}=(X_{i}-X_{j})^{T}\beta+(z_{i}-z_{j})$. The new covariate $X_{i}-X_{j}$ certainly satisfies symmetry regardless of the distribution of $X_{i}$.
We finally give a brief discussion on the condition $\frac{p^{2}}{n(1-\epsilon)^{4}}\leq c$ for some sufficiently small constant $c>0$. When $\epsilon$ is bounded away from $1$, this condition allows $p$ to grow with $n$ as long as $n\gg p^{2}$. The requirement that the sample size needs to exceed the squared degrees of freedom has been previously noted in the high-dimensional literature when the signal is very sparse (Zhang and Zhang,, 2014; Javanmard and Montanari,, 2014; van de Geer et al.,, 2014), and can be shown to be necessary for interval estimation with parametric rate (Ren et al.,, 2015; Cai and Guo,, 2017; Verzelen and Gassiat,, 2018). When the signal is dense, however, the condition $n\gg p^{2}$ is not necessary for confidence interval construction in the absence of contamination (Portnoy,, 1985; Wen et al.,, 2025). In contrast, under Huber’s contamination model (2), the optimal sample complexity for confidence interval construction remains unknown. In particular, it is unclear whether the quadratic dependence on $p$ is information-theoretically necessary, even without computational constraints, or whether it can be improved by alternative procedures. Establishing the exact dependence on $p$ and $\epsilon$, as well as understanding possible gaps between information-theoretic and computational limits, are important open problems for future work.
@@ 3.1 A Symmetric Location Model
where $z_{i}\sim(1-\epsilon)N(0,1)+\epsilon Q_{0}$ for some distribution $Q_{0}$ that is symmetric around zero. In other words, $z_{i}\sim Q_{0}$ implies $z_{i}\sim Q_{0}$. We emphasize that such a symmetry assumption is not assumed for the error distribution in (9).
@@ 3.3 Application in Robust Regression
Since the covariate $x_{i}$ follows $N(0,1)$ and the noise variable $z_{i}$ is independent from $x_{i}$, the random variable $\frac{z_{i}}{x_{i}}$ is symmetric around zero regardless of the noise distribution. Therefore, (15) is an instance of the symmetric location model (9), where $\beta$ is the center of the distribution. Thus, the regression coefficient $\beta$ is identifiable even when $\epsilon=1$. When the distribution of $x_{i}$ is not necessarily $N(0,1)$ or not even symmetric, one could still deal with a linear model with symmetric design by preprocessing the data with pairwise difference $y_{i}-y_{j}=\beta(x_{i}-x_{j})+z_{i}-z_{j}$.
@@ 4.1 Decorrelating Nuisance Parameters
Even though $\widetilde{x}_{i}$ and $(\gamma-\widehat{\gamma})^{T}w_{i}+z_{i}$ are still not independent in (27), the independence between $\widetilde{x}_{i}$ and $z_{i}$ and the uncorrelatedness between $\widetilde{x}_{i}$ and $w_{i}$ are sufficient for us to apply (21). We emphasize that the uncorrelatedness condition (28) is only in the sample level, which turns out to be particularly convenient when bounding the model misspecification error. Moreover, the $\widehat{\gamma}$ in (26) is a robust estimator for $\gamma=\beta\widehat{\alpha}+\theta$, which will ensure that the scale of the noise variable $(\gamma-\widehat{\gamma})^{T}w_{i}+z_{i}$ is well controlled.
@@ 4.2 Coverage
To analyze the effect of model misspecification caused by $(\gamma-\widehat{\gamma})^{T}w_{i}$, we use the smoothness of $g(\cdot)$ and have
@@ 4.3 Length
4.4 Linear Functional
Algorithm 1 can also be applied to robust confidence interval construction of a general linear functional $u^{T}b$ in the linear model (1). The only modification is an additional rotation of the design matrix before applying Algorithm 1. Let $V=(v_{1},\cdots,v_{p})\in\mathbb{R}^{p\times p}$ be an orthonormal matrix with the first column given by $v_{1}=u/\|u\|$. Then, the regression model $X_{i}^{T}b$ can be written as
$X_{i}^{T}b=X_{i}^{T}\left(VV^{T}\right)b=\frac{1}{\|u\|^{2}}X_{i}^{T}uu^{T}b+\sum_{j=2}^{p}X_{i}^{T}v_{j}v_{j}^{T}b.$
Then, the output of Algorithm 1 with the same $y_{i}$ and the new covariates
$\left(\frac{1}{\|u\|^{2}}X_{i}^{T}u,X_{i}^{T}v_{2},\cdots,X_{i}^{T}v_{p}\right)^{T}$
is a valid robust confidence interval of $u^{T}b$.
@@ 5.1 Numerical Setups
    Gaussian: $z_{i}\overset{iid}{\sim}(1-\epsilon)N(0,1)+\epsilon\left(\frac{1}{2}N(10^{2},10^{4})+\frac{1}{2}N(10^{4},10^{8})\right)$.
@@
    Cauchy: $z_{i}\overset{iid}{\sim}(1-\epsilon)C(0,1)+\epsilon\left(\frac{1}{2}N(10^{2},10^{4})+\frac{1}{2}N(10^{4},10^{8})\right)$.
@@
The threshold level $1.5\sqrt{\log(2/\alpha)}$ used in Algorithm 1 is to guarantee non-asymptotic coverage in Theorem 1. One may wonder whether replacing $1.5\sqrt{\log(2/\alpha)}$ with the more aggressive Gaussian quantile $\Phi^{-1}(1-\alpha/2)$ works better in practice. While the Gaussian approximation works in the asymptotics of $\frac{p^{2}}{n(1-\epsilon)^{4}}\rightarrow 0$, our numerical results indicate that it should be used with caution when the sample size $n$ is small or $\epsilon$ is large. In addition to this variation of Algorithm 1, we also implement several methods in the literature. This includes the residual bootstrap (RB) of Freedman, (1981), the restructured regression (RR) of Zhu and Bradic, (2018), the cyclic permutation test (CPT) of Lei and Bickel, (2021), the convex hull method (HulC) of Kuchibhotla et al., (2024) and the residual permutation test (RPT) of Wen et al., (2025). We also report the performance of the classical confidence interval computed from the t-statistic centering at the ordinary least-squares estimator (OLS). Among the above methods, RR (Zhu and Bradic,, 2018), CPT (Lei and Bickel,, 2021) and RPT (Wen et al.,, 2025) are originally proposed to test the null hypothesis that a coefficient is zero, which also lead to confidence intervals by inverting the tests.
@@ 5.2 Implementation Details
where $\rho(\cdot)$ is the same Huber loss used in Algorithm 1. The bootstrap sample size is taken as $200$. The restructured regression method has two versions in Zhu and Bradic, (2018), and we implement the one with the knowledge of the design covariance. The cyclic permutation test (Lei and Bickel,, 2021) is implemented with the recommended preordering computed using a genetic algorithm. The genetic algorithm is computed with $1000$ random samples. The HulC method (Kuchibhotla et al.,, 2024) is also implemented with the Huber regression (45), together with a recommended randomized choice of sample splitting to prevent the interval from being overly conservative. The residual permutation test (Wen et al.,, 2025) is implemented with the more powerful version using the empirical p-values, and the tuning parameters are set as recommended in the original work.
@@ 5.3 Results and Discussion
5.4 Sensitivity to Model Parameters
We conduct additional simulations to assess the sensitivity of the proposed method to key parameters, including sample size, scaling parameter, and the covariate distribution. The observations are generated according to the same setup described in Section 5.1 unless otherwise mentioned. The noise distribution is Gaussian and the confidence level is set as $1-\alpha=0.95$. All experiments are implemented by Alg1 (the original version of Algorithm 1). Since we observe that the coverage property is always satisfied, we only report the length.
Figure 3: Average interval length against model parameters and covariate distribution.
Figure 3(a) varies $n\in[40,200]$ with $p=20$ and $\epsilon=0.2$ being fixed. The goal is to check how soon the robustness property of Algorithm 1 kicks in as the sample size increases. For this setting with moderate dimension, $n\geq 50$ would be sufficient.
We use Figure 3(b) to illustrate the effect of the algorithm scaling. Suppose we do not know that $\sigma^{2}=1$, Algorithm 1 should be applied with the scaling $g(\cdot)\rightarrow g(\cdot/\tau)$ and $\rho(\cdot)\rightarrow\rho(\cdot/\tau)$ for some $\tau>0$. Then, Theorem 1 continues to hold as long as $\tau\lesssim 1$ and $\frac{p^{2}}{n(1-\epsilon)^{4}\tau^{4}}$ is sufficiently small. Figure 3(b) shows that this is indeed the case. With $n=1000$, $p=20$ and $\epsilon=0.2$ being fixed, the length of the confidence interval stabilizes as long as $\tau$ is set to be sufficiently small. In fact, there does not seem to exist a regime where $\tau$ is too small, despite that we still need the condition $\frac{p^{2}}{n(1-\epsilon)^{4}\tau^{4}}<c$ theoretically.
Figure 3(c) investigates the effect of the covariate distribution. In particular, we generate $X_{i}$ so that $\Sigma^{-1/2}X_{i}$ consists of $p$ independent standard Cauchy, $t$-distribution (3 degrees of freedom), Laplace and Gaussian with $\Sigma_{jk}=0.6^{|j-k|}$. The simulation is conducted with $p=20$ and $n=1000$ with $\epsilon$ varying from 0 to $0.8$. It is clear from Figure 3(c) that a lighter tail tends to be more robust, which is naturally expected. On the other hand, the result is not sensitive to the distribution tail. The breakdown point of the case with $t$-distribution still exceeds $\epsilon=0.7$.
Figure 4: Scaled average interval length against $\epsilon$ with various correlation structures.
We finally investigate the effect of the correlation structure of the covariates. Let $X_{i}\overset{iid}{\sim}N(0,\Sigma)$. We consider three types of covariance matrices: equicorrelation, $\Sigma_{jk}=\rho+(1-\rho){\mathds{1}}\{j=k\}$; Toeplitz, $\Sigma_{jk}=\rho^{|j-k|}$; and block Toeplitz, where $\rho^{|j-k|}$ within each of five equal-sized independent blocks. For a given $\Sigma$, the effective noise level for inference on the first regression coefficient is $\sqrt{(\Sigma^{-1})_{11}}$. Accordingly, we report the length of the confidence interval scaled by the inverse of this noise level. The simulation results for $p=20$ and $n=1000$ are shown in Figure 4. The curves of the scaled confidence interval length as functions of $\epsilon$ are well aligned across different values of $\rho$ and across the three covariance structures.
5.5 A Real Data Example
We apply our robust algorithm to the Boston Housing data. The data were originally collected for the study of Harrison and Rubinfeld, (1978) on housing prices. The variable description and linear model can be found at Table IV and the appendix of Harrison and Rubinfeld, (1978). The variables are transformed and standardized before regression.
For each regression coefficient, instead of computing its confidence interval, we modify Algorithm 1 to compute a point estimator and a $p$-value. With $(\widetilde{x_{i}},\widetilde{y}_{i})$ computed by the first four steps of Algorithm 1, the point estimator is found as the solution to $\sum_{i=1}^{n}g(\widetilde{y}_{i}-\widetilde{x}_{i}\beta)\widetilde{x}_{i}=0$, and the $p$-value is given by the formula
$2{\rm{exp}}\left(-\frac{\left|\sum_{i=1}^{n}g(\widetilde{y}_{i})\widetilde{x}_{i}\right|^{2}}{2.25\sum_{i=1}^{n}\widetilde{x}_{i}^{2}}\right).$
We first apply OLS and Algorithm 1 to the original data. Then, Algorithm 1 is applied again to the contaminated data after adding $\frac{1}{2}N(10^{2},10^{4})+\frac{1}{2}N(10^{4},10^{8})$ to each response variable with probability $0.1$. The estimated coefficients and $p$-values are reported in Table 1.
Table 1: Estimated coefficients and $p$-values for the Boston Housing data.
OLS Alg1 Alg1_c
Covariate Coef. $p$-value Coef. $p$-value Coef. $p$-value
age 0.006 0.863 -0.039 1 -0.039 1
chas 0.057 0.006 0.006 1 0.001 1
crim -0.250 <0.001 -0.268 <0.001 -0.318 0.023
dis -0.252 <0.001 -0.283 0.002 -0.228 0.137
indus 0.004 0.919 -0.053 0.805 0.120 0.829
nox -0.217 <0.001 -0.169 0.007 -0.143 0.306
ptratio -0.165 <0.001 -0.137 <0.001 -0.182 <0.001
rad 0.205 <0.001 0.135 0.013 0.234 0.037
rm 0.141 <0.001 0.203 0.001 0.281 0.043
tax -0.173 <0.001 -0.138 0.228 -0.300 0.069
zn 0.005 0.874 -0.004 1 0.042 1
According to the table, the estimated coefficients are relatively stable in terms of magnitude and sign. The inference, however, shows more variability. The OLS identifies many covariates as significant, while Algorithm 1 is more conservative and identifies fewer significant covariates. Since the residual Q–Q plot is quite heavy-tailed, the results of OLS are less reliable. Moreover, after the data is contaminated, Algorithm 1 selects even fewer covariates as significant. The covariates crim, ptratio, rad and rm are significant under all scenarios, revealing a strong dependence of housing price on crime rate, pupil-teacher ratio, accessibility to radial highways, and average room number.
6 Beyond Oblivious Contamination
The robust regression setting (1) with noise variables (2) has two crucial conditions: first, the contamination does not affect the covariates $X_{i}$; second, the contamination is independent of $X_{i}$. While relaxing these two conditions is certainly possible for estimation, we show in this section that the two conditions are essentially necessary for adaptive inference.
Joint Contamination on $(X_{i},y_{i})$ Let us consider the univariate setting where samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ are i.i.d. generated from the Huber contamination model $(1-\epsilon)P_{\beta,\tau,\sigma}+\epsilon Q$. Here, $P_{\beta,\tau,\sigma}$ stands for the joint distribution of $(x_{i},y_{i})$ whose sampling process is $x_{i}\sim N(0,\tau^{2})$ and $y_{i}|x_{i}\sim N(\beta x_{i},\sigma^{2})$. In this setting where both $x_{i}$ and $y_{i}$ are contaminated, a shrinking adaptive confidence interval of $\beta$ is impossible.
Proposition 3.
There exists some $Q$ and some sufficiently small constant $c$, such that $P_{0,1,1}=\frac{2}{3}P_{c,0.99,0.99}+\frac{1}{3}Q$.
The proposition shows that a linear model with $\beta=0$ and no contamination can be represented as a linear model with $\beta=c$ under $1/3$ contamination. Here, the constant $c$ can be taken as $\frac{1-0.99^{2}}{2\times 0.99}$. Therefore, any confidence interval that is agnostic between $\epsilon=0$ and $\epsilon=1/3$ must contain both 0 and $c$, and hence has a constant length independent of the sample size.
Contamination Depending on $X_{i}$ Even if the contamination only affects the response variables $y_{i}$, it will still be impossible to construct a shrinking adaptive confidence interval unless the contamination is independent of $X_{i}$. Let us consider the univariate setting, and the relation of the pair $(x_{i},y_{i})$ follows
$y_{i}=\beta x_{i}+z_{i}+o_{i},$ (46)
where $\beta\in\mathbb{R}$ and $o_{i}$ is an outlier variable that may depend on $x_{i}$. The model (46) has been widely studied in the literature on robust regression Gannaz, (2007); McCann and Welsch, (2007); She and Owen, (2011). When $o_{i}$ is independent of $x_{i}$, the sum $z_{i}+o_{i}$ follows (2) and is thus a special case of the robust regression with oblivious contamination.
We introduce the $P_{\beta,\sigma,Q(\cdot)}$ as the joint distribution of $(x_{i},y_{i})$, whose sampling process is given by the linear equation (46), where $x_{i}\sim N(0,1)$ and then $z_{i}\sim N(0,\sigma^{2})$ independently of $x_{i}$. Finally, $o_{i}|x_{i},z_{i}\sim Q(x_{i},z_{i})$ for the conditional distribution $Q(\cdot)$ given $x_{i}$ and $z_{i}$. The contamination proportion is $\epsilon=\mathbb{E}Q(o_{i}\neq 0|x_{i},z_{i})$.
Proposition 4.
There exist $Q_{0}(\cdot)$, $Q_{1}(\cdot)$ and some sufficiently small constant $c$, such that $\mathbb{E}Q_{0}(o_{i}\neq 0|x_{i},z_{i})=0$, $\mathbb{E}Q_{1}(o_{i}\neq 0|x_{i},z_{i})=0.4$, and $P_{0,1,Q_{0}(\cdot)}=P_{c,0.99,Q_{1}(\cdot)}$.
Proposition 4 admits an interpretation similar to that of Proposition 3. In particular, a linear model with $\beta=0$ and no contamination can be represented as a linear model with $\beta=c$ under $40\%$ contamination where $c$ can be taken as 0.07. Consequently, any confidence interval that is agnostic about the contamination proportion must have a constant length independent of the sample size.